OpenStudy (anonymous):

cos^2(x)+cos(x)=cos(2x)

OpenStudy (anonymous):

find x ?

OpenStudy (anonymous):

It's asking to find all solutions in interval \[[0,2\pi)\]

OpenStudy (anonymous):

cos^2x - cos x = 0 cos x ( cos 2x - 1) = 0 cos x = 0 or cos 2x -1 = 0 cos 2x =1 cos x =0 , cos 2x = 1 see if i did wrong in my steps ...

OpenStudy (anonymous):

x = cos-1 ( 0) 2x = cos-1 (1) x = Pi/2

OpenStudy (anonymous):

2nd value : x = 2Pi - Pi/2 = 3/2 Pi

OpenStudy (anonymous):

cos2x=2cos^2x-1 would be the first step I would use. Then move everything to one side and get: cos^2x-cosx-1=0 From there, it doesn't factor so you will have to use a quadratic fromula.